# Gradient-Based Multi-Objective Deep Learning: Algorithms, Theories, Applications, and Beyond

### v2

@@ 5.2 Training Strategy
$\min_{\bm{\phi}}\mathbb{E}_{{\bm{\alpha}}\sim\Delta_{m-1}}~{}\mathbb{E}_{({\bm{x}},{\bm{y}})\sim\mathcal{D}}\left[\tilde{g}_{\bm{\alpha}}({\bm{\ell}}(t({\bm{x}};{\bm{\theta}}({\bm{\alpha}};{\bm{\phi}}),{\bm{y}}))\right],$ (40)
where $({\bm{x}},{\bm{y}})$’s are training data sampled from dataset $\mathcal{D}$, ${\bm{\ell}}(\cdot,\cdot)$ is the multi-objective loss function and $\tilde{g}_{\bm{\alpha}}(\cdot)$ is an MOO algorithm that produces a single solution given preference vector ${\bm{\alpha}}$. In general, most optimization algorithms discussed in Sections 2.2 and 4 can be adopted. Below, we summarize the algorithms adopted by the existing methods.
  • Scalarization: As outlined in Section 2.2, scalarization combines multiple objectives into one scalar objective. Linear scalarization is common in most algorithms, such as PHN [52], COSMOS [57], PAMAL [60], and LORPMAN [61]. Alternatives like Tchebycheff and smooth Tchebycheff scalarization [18] can also be adopted.
  • Preference-Aware Weighting Methods: As discussed in Section 4.1, these methods incorporate user preferences into optimization. PHN [52] employs the Exact Pareto Optimal (EPO) solver [37] to find Pareto-optimal solutions aligned with preferences. CPMTL [53] uses a constrained approach inspired by PMTL [36].
  • Hypervolume Maximization: As discussed in Section 4.2, hypervolume maximization can optimize both diversity and convergence of a solution set. PHN-HVI [54] leverages hypervolume maximization to optimize the structural parameters ${\bm{\phi}}$. In each iteration, it samples a set of preference vectors and generates solutions based on these vectors. Subsequently, hypervolume maximization is applied to promote both the diversity and convergence of the solutions.
5.3 Discussions
@@
6 Theories
6.1 Convergence of Gradient-Balancing Methods in Section 3.2
This section reviews the convergence analysis for gradient-balancing methods in Section 3.2. This is first examined under the deterministic gradient setting (i.e., full-batch gradient). Subsequently, it is analyzed under the stochastic gradient setting. Note that in non-convex MOO, convergence refers to reaching Pareto stationary (Definition 5).
6.1.1 Deterministic Gradient
With the use of deterministic gradients, it has shown that under mild conditions, MGDA [6] can converge to a Pareto-stationary point at a convergence rate of $\mathcal{O}(K^{-1/2})$ [97], where $K$ is the number of iterations. This rate is comparable to that of single-objective optimization [98]. The fundamental idea is to assess the reduction in each individual objective function. When the current solution is not Pareto-stationary, it can be shown that the common descent direction ${\bm{d}}$ identified by MGDA ensures an improvement of at least $\mathcal{O}(\left\|{\bm{d}}\right\|^{2})$ in each objective, which mirrors the situation in single-objective optimization [97]. Therefore, the method used in single-objective optimization can be applied to prove the convergence. Furthermore, other algorithms such as CAGrad [19], Nash-MTL [27], and PCGrad [33] also offer convergence analyses in similar approaches.
TABLE II: Convergence of existing stochastic gradient MOO algorithms. “LS” and “GS” denote the objectives are $L$-smooth and generalized $L$-smooth, respectively. “BG” and “BF” represent the bounded gradient and bounded function value assumptions, respectively. “Complexity” denotes the sample complexity achieving $\epsilon$-accurate Pareto stationary point. “Bounded CA” denotes the bounded conflict-avoidant distance.
@@ 6.1.1 Deterministic Gradient
SMG [20] ${\mathcal{O}}(\epsilon^{-2})$ LS, BG ${\mathcal{O}}(\epsilon^{-4})$
CR-MOGM [21] ${\mathcal{O}}(1)$ LS, BF, BG ${\mathcal{O}}(\epsilon^{-2})$
MoCo [22] ${\mathcal{O}}(1)$ LS, BF, BG ${\mathcal{O}}(\epsilon^{-2})$
PSMGD [23] ${\mathcal{O}}(1)$ LS, BF, BG ${\mathcal{O}}(\epsilon^{-2})$
MoDo [24] ${\mathcal{O}}(1)$ LS, BG ${\mathcal{O}}(\epsilon^{-2})$
SDMGrad [25] ${\mathcal{O}}(1)$ LS, BG ${\mathcal{O}}(\epsilon^{-2})$
SGSMGrad [26] ${\mathcal{O}}(1)$ GS ${\mathcal{O}}(\epsilon^{-2})$
6.1.2 Stochastic Gradient
Liu and Vicente [20] are the first to analyze the convergence of gradient-based MOO algorithms with stochastic gradients. They propose a stochastic version of MGDA [6] that computes a common descent direction using stochastic gradients for all objectives and prove convergence to Pareto-optimal solutions under the assumption of convex objective functions. However, due to the inherent bias of common descent direction ${\bm{d}}$ introduced by stochastic gradient estimations, their analysis requires the use of an increasing batch size that grows linearly with the number of iterations to ensure convergence. This requirement can be impractical in real-world applications, as large batch sizes lead to increased computational costs and memory usage.
To overcome this limitation, Zhou et al. [21] propose Correlation-Reduced Stochastic Multi-Objective Gradient Manipulation (CR-MOGM). It addresses the bias in the common descent direction by introducing a smoothing technique on weight coefficient ${\bm{\lambda}}$. In CR-MOGM, ${\bm{\lambda}}^{(k)}$ at iteration $k$ is updated using a moving average:
${\bm{\lambda}}^{(k)}=(1-\gamma)\hat{{\bm{\lambda}}}^{(k)}+\gamma{\bm{\lambda}}^{(k-1)},$ (41)
where $\gamma\in[0,1]$ is a smoothing factor, $\hat{{\bm{\lambda}}}^{(k)}$ is the weight vector obtained by the MOO solver (such as MGDA) at iteration $k$ using the current stochastic gradients, and ${\bm{\lambda}}^{(k-1)}$ is the smoothed weight from the previous iteration. Smoothing reduces the variance in the weight, leading to a more stable and reliable common descent direction. Xu et al. [23] further analyze the convergence by updating the objective weights every $\tau$ iterations instead of at each iteration, and achieve the same convergence rate ${\mathcal{O}}(\epsilon^{-2})$ as CR-MOGM. They also match the convergence rate in stochastic single-objective optimization [99].
Another approach to address gradient bias is MoCo [22], which introduces a tracking variable $\hat{{\bm{g}}}_{i}^{(k)}$ that approximates the true gradient:
$\hat{{\bm{g}}}^{(k+1)}_{i}=\prod_{L_{i}}\left(\hat{{\bm{g}}}^{(k)}_{i}-\gamma\left(\hat{{\bm{g}}}^{(k)}_{i}-{\bm{g}}^{(k)}_{i}\right)\right),$ (42)
@@ 6.1.2 Stochastic Gradient
To address these limitations, Chen et al. [24] propose the Multi-objective gradient with double sampling algorithm (MoDo). MoDo mitigates bias in stochastic gradient-based MOO methods without assuming bounded function values. It introduces a double sampling technique to obtain unbiased estimates of the gradient products needed for updating the weight coefficients. Specifically, at each iteration, MoDo updates the weight vector ${\bm{\lambda}}^{(k)}$ using gradients computed on two independent mini-batches:
${\bm{\lambda}}^{(k+1)}=\prod_{\Delta_{m-1}}\left({\bm{\lambda}}^{(k)}-\eta{\bm{G}}^{(k)}({{\bm{z}}_{1}^{(k)}})^{\top}{\bm{G}}^{(k)}({{\bm{z}}^{(k)}_{2}}){\bm{\lambda}}^{(k)}\right),$ (43)
@@
Similarly, Xiao et al. [25] propose the Stochastic Direction-oriented Multi-objective Gradient descent (SDMGrad) algorithm, which introduces a new direction-oriented multi-objective formulation by regularizing the common descent direction within a neighborhood of a target direction (such as the average gradient of all objectives). In SDMGrad, the weight vector ${\bm{\lambda}}^{(k)}$ is updated as:
${\bm{\lambda}}^{(k+1)}=\prod_{\Delta_{m-1}}\left({\bm{\lambda}}^{(k)}-\eta\left[{\bm{G}}^{(k)}({{\bm{z}}^{(k)}_{1}})^{\top}\left({\bm{G}}^{(k)}({\bm{z}}_{2}^{(k)}){\bm{\lambda}}^{(k)}\right.\right.\right.$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\left.\left.\left.+\gamma{\bm{g}}_{0}({\bm{z}}^{(k)}_{2})\right)\right]\right),$ (44)
where $\eta$ is the step size, $\gamma$ is a regularization factor controlling the proximity to the target direction ${\bm{g}}_{0}({\bm{z}}_{2}^{(k)})$. By incorporating this regularization, SDMGrad effectively balances optimization across objectives while guiding the overall descent direction. Zhang et al. [26] further consider the convergence under generalized $L$-smoothness [100] and without bounded gradient assumption. A summary of the convergence results of existing stochastic gradient MOO algorithms is provided in Table II.
6.2 Generalization
The generalization aspect of multi-objective deep learning remains relatively underexplored compared to its convergence properties. Cortes et al. [101] analyze the generalization behavior of a specific scalarization approach that minimizes over convex combinations. Sukenik et al. [102] consider a broader class of scalarization methods. They show that the generalization bounds for individual objectives extend to MOO with scalarization. Both studies rely on the Rademacher complexity of the hypothesis class to establish algorithm-independent generalization bounds, which are unaffected by the training process.
@@ 6.2 Generalization
$\max_{i\in[m]}\ell_{{\mathcal{D}}_{i}}(h)\leq\min_{h^{*}\in{\mathcal{H}}}\max_{i\in[m]}\ell_{{\mathcal{D}}_{i}}(h^{*})+\epsilon,$ (45)
where $\ell_{{\mathcal{D}}_{i}}(h)$ is the loss of hypothesis $h$ on ${\mathcal{D}}_{i}$, $h^{*}$ is the optimal hypothesis and $\epsilon$ is the error tolerance. It is first formulated by Awasthi et al. [103] as an open problem in 2023 and has since been addressed by several works [104, 105]. Haghtalab et al. are the first to show the sample complexity lower bound of $\widetilde{\Omega}\left(\frac{v+m}{\epsilon^{2}}\right)$, where $v=\operatorname{VCdim}({\mathcal{H}})$ is the VC-dimension of the hypothesis class ${\mathcal{H}}$. By using a boosting framework, Peng [104] gives an algorithm that achieves a sample complexity upper bound of $\widetilde{O}\left(\frac{v+m}{\epsilon^{2}}\cdot\left(\frac{m}{\epsilon}\right)^{o(1)}\right)$, which is nearly optimal. In a concurrent work, Zhang et al. [105] provide a variant of the hedge algorithm under an Empirical Risk Minimization (ERM) oracle access that achieves the optimal sample complexity bound of $\widetilde{O}\left(\frac{v+m}{\epsilon^{2}}\right)$.
In the online setting, Liu et al. [106] provide an adaptive online mirror descent algorithm that achieves $\mathcal{O}\left(\frac{mv}{\sqrt{K}}\right)$ regret, where $K$ is the number of iterations. Using a plain online-to-batch conversion scheme, this algorithm leads to an $\mathcal{O}\left(\frac{m^{2}v^{2}}{\epsilon^{2}}\right)$ sample complexity, which still lags behind the optimal offline sample complexity of $\widetilde{O}\left(\frac{m+v}{\epsilon^{2}}\right)$ [105]. It is interesting to see whether it is possible to use an online learner with proper online-to-batch conversion schemes that is able to match the optimal offline sample complexity.
MoDo [24] introduces a different perspective by leveraging algorithm stability to derive algorithm-dependent generalization error bounds of gradient balancing algorithms.
7 Applications
7.1 Reinforcement Learning
Reinforcement Learning (RL) [107] involves training agents to make sequential decisions by maximizing cumulative rewards. Multi-Objective Reinforcement Learning (MORL) [108] extends single-objective RL by utilizing a vector-valued reward function ${\bm{r}}(s,a):\mathcal{S}\times\mathcal{A}\mapsto\mathbb{R}^{m}$. MORL aims to learn a policy network $\pi_{\bm{\theta}}(s)$ with the following training objective:
$\min_{{\bm{\theta}}}{\bm{f}}({\bm{\theta}}):=\mathbb{E}_{a\sim\pi_{\bm{\theta}}(s)}\sum_{t=1}^{\infty}\gamma^{t}{\bm{r}}(s_{t},a_{t}),$ (46)

### v3

@@ 5.2. Training Strategy
(39) $\min_{\bm{\phi}}\mathbb{E}_{{\bm{\alpha}}\sim\Delta_{m-1}}~\mathbb{E}_{({\bm{x}},{\bm{y}})\sim\mathcal{D}}\left[\tilde{g}_{\bm{\alpha}}({\bm{\ell}}(t({\bm{x}};{\bm{\theta}}({\bm{\alpha}};{\bm{\phi}}),{\bm{y}}))\right],$
[…]
  • Scalarization: As outlined in Section 2.2, scalarization combines multiple objectives into one scalar objective. Linear scalarization is common in most algorithms, such as PHN (Navon et al., 2021), COSMOS (Ruchte and Grabocka, 2021), PAMAL (Dimitriadis et al., 2023), and LORPMAN (Chen and Kwok, 2024a). Alternatives like Tchebycheff and smooth Tchebycheff scalarization (Lin et al., 2024b) can also be adopted.
  • Preference-Aware Weighting Methods: As discussed in Section 4.1, these methods incorporate user preferences into optimization. PHN (Navon et al., 2021) employs the Exact Pareto Optimal (EPO) solver (Mahapatra and Rajan, 2020) to find Pareto-optimal solutions aligned with preferences. CPMTL (Lin et al., 2020) uses a constrained approach inspired by PMTL (Lin et al., 2019b).
  • Hypervolume Maximization: As discussed in Section 4.2, hypervolume maximization can optimize both diversity and convergence of a solution set. PHN-HVI (Hoang et al., 2023) leverages hypervolume maximization to optimize the structural parameters $\bm{\phi}$}. In each iteration, it samples a set of preference vectors and generates solutions based on these vectors. Subsequently, hypervolume maximization is applied to promote both the diversity and convergence of the solutions.
5.3. Discussions
@@
6. Theories
6.1. Convergence of Gradient-Balancing Methods in Section 3.2
This section reviews the convergence analysis for gradient-balancing methods in Section 3.2. This is first examined under the deterministic gradient setting (i.e., full-batch gradient). Subsequently, it is analyzed under the stochastic gradient setting. Note that in non-convex MOO, convergence refers to reaching Pareto stationary (Definition 2.5).
6.1.1. Deterministic Gradient
With the use of deterministic gradients, it has shown that under mild conditions, MGDA (Sener and Koltun, 2018) can converge to a Pareto-stationary point at a convergence rate of $\mathcal{O}(K^{-1/2})$ (Fliege et al., 2019), where $K$ is the number of iterations. This rate is comparable to that of single-objective optimization (Nesterov, 2013). The fundamental idea is to assess the reduction in each individual objective function. When the current solution is not Pareto-stationary, it can be shown that the common descent direction $\bm{d}$} identified by MGDA ensures an improvement of at least $\mathcal{O}(\left\|{\bm{d}}\right\|^{2})$ in each objective, which mirrors the situation in single-objective optimization (Fliege et al., 2019). Therefore, the method used in single-objective optimization can be applied to prove the convergence. Furthermore, other algorithms such as CAGrad (Liu et al., 2021b), Nash-MTL (Navon et al., 2022), and PCGrad (Yu et al., 2020a) also offer convergence analyses in similar approaches.
6.1.2. Stochastic Gradient
Liu and Vicente (Liu and Vicente, 2021) are the first to analyze the convergence of gradient-based MOO algorithms with stochastic gradients. They propose a stochastic version of MGDA (Sener and Koltun, 2018) that computes a common descent direction using stochastic gradients for all objectives and prove convergence to Pareto-optimal solutions under the assumption of convex objective functions. However, due to the inherent bias of common descent direction $\bm{d}$} introduced by stochastic gradient estimations, their analysis requires the use of an increasing batch size that grows linearly with the number of iterations to ensure convergence. This requirement can be impractical in real-world applications, as large batch sizes lead to increased computational costs and memory usage.
Table 2. Convergence of existing stochastic gradient MOO algorithms. “LS” and “GS” denote the objectives are $L$-smooth and generalized $L$-smooth, respectively. “BG” and “BF” represent the bounded gradient and bounded function value assumptions, respectively. “Complexity” denotes the sample complexity achieving $\epsilon$-accurate Pareto stationary point. “Bounded CA” denotes the bounded conflict-avoidant distance.
@@ 6.1.2. Stochastic Gradient
SMG (Liu and Vicente, 2021) ${\mathcal{O}}(\epsilon^{-2})$ LS, BG ${\mathcal{O}}(\epsilon^{-4})$
CR-MOGM (Zhou et al., 2022) ${\mathcal{O}}(1)$ LS, BF, BG ${\mathcal{O}}(\epsilon^{-2})$
MoCo (Fernando et al., 2023) ${\mathcal{O}}(1)$ LS, BF, BG ${\mathcal{O}}(\epsilon^{-2})$
PSMGD (Xu et al., 2025b) ${\mathcal{O}}(1)$ LS, BF, BG ${\mathcal{O}}(\epsilon^{-2})$
MoDo (Chen et al., 2023a) ${\mathcal{O}}(1)$ LS, BG ${\mathcal{O}}(\epsilon^{-2})$
SDMGrad (Xiao et al., 2023) ${\mathcal{O}}(1)$ LS, BG ${\mathcal{O}}(\epsilon^{-2})$
SGSMGrad (Zhang et al., 2025) ${\mathcal{O}}(1)$ GS ${\mathcal{O}}(\epsilon^{-2})$
To overcome this limitation, Zhou et al. (Zhou et al., 2022) propose Correlation-Reduced Stochastic Multi-Objective Gradient Manipulation (CR-MOGM). It addresses the bias in the common descent direction by introducing a smoothing technique on weight coefficient $\bm{\lambda}$}. In CR-MOGM, ${\bm{\lambda}}^{(k)}$ at iteration $k$ is updated using a moving average:
(40) ${\bm{\lambda}}^{(k)}=(1-\gamma)\hat{{\bm{\lambda}}}^{(k)}+\gamma{\bm{\lambda}}^{(k-1)},$
where $\gamma\in[0,1]$ is a smoothing factor, $\hat{{\bm{\lambda}}}^{(k)}$ is the weight vector obtained by the MOO solver (such as MGDA) at iteration $k$ using the current stochastic gradients, and ${\bm{\lambda}}^{(k-1)}$ is the smoothed weight from the previous iteration. Smoothing reduces the variance in the weight, leading to a more stable and reliable common descent direction. Xu et al. (Xu et al., 2025b) further analyze the convergence by updating the objective weights every $\tau$ iterations instead of at each iteration, and achieve the same convergence rate ${\mathcal{O}}(\epsilon^{-2})$ as CR-MOGM. They also match the convergence rate in stochastic single-objective optimization (Ghadimi and Lan, 2013).
Another approach to address gradient bias is MoCo (Fernando et al., 2023), which introduces a tracking variable $\hat{{\bm{g}}}_{i}^{(k)}$ that approximates the true gradient:
(41) $\hat{{\bm{g}}}^{(k+1)}_{i}=\prod_{L_{i}}\left(\hat{{\bm{g}}}^{(k)}_{i}-\gamma\left(\hat{{\bm{g}}}^{(k)}_{i}-{\bm{g}}^{(k)}_{i}\right)\right)$,
@@
To address these limitations, Chen et al. (Chen et al., 2023a) propose the Multi-objective gradient with double sampling algorithm (MoDo). MoDo mitigates bias in stochastic gradient-based MOO methods without assuming bounded function values. It introduces a double sampling technique to obtain unbiased estimates of the gradient products needed for updating the weight coefficients. Specifically, at each iteration, MoDo updates the weight vector ${\bm{\lambda}}^{(k)}$ using gradients computed on two independent mini-batches:
(42) ${\bm{\lambda}}^{(k+1)}=\prod_{\Delta_{m-1}}\left({\bm{\lambda}}^{(k)}-\eta{\bm{G}}^{(k)}({{\bm{z}}_{1}^{(k)}})^{\top}{\bm{G}}^{(k)}({{\bm{z}}^{(k)}_{2}}){\bm{\lambda}}^{(k)}\right),$
@@
Similarly, Xiao et al. (Xiao et al., 2023) propose the Stochastic Direction-oriented Multi-objective Gradient descent (SDMGrad) algorithm, which introduces a new direction-oriented multi-objective formulation by regularizing the common descent direction within a neighborhood of a target direction (such as the average gradient of all objectives). In SDMGrad, the weight vector ${\bm{\lambda}}^{(k)}$ is updated as:
[…]
where $\eta$ is the step size, $\gamma$ is a regularization factor controlling the proximity to the target direction ${\bm{g}}_{0}({\bm{z}}_{2}^{(k)})$. By incorporating this regularization, SDMGrad effectively balances optimization across objectives while guiding the overall descent direction. Zhang et al. (Zhang et al., 2025) further consider the convergence under generalized $L$-smoothness (Li et al., 2023) and without bounded gradient assumption. A summary of the convergence results of existing stochastic gradient MOO algorithms is provided in Table 2.
6.2. Generalization
The generalization aspect of multi-objective deep learning remains relatively underexplored compared to its convergence properties. Cortes et al. (Cortes et al., 2020) analyze the generalization behavior of a specific scalarization approach that minimizes over convex combinations. Sukenik et al. (Súken´ık and Lampert, 2024) consider a broader class of scalarization methods. They show that the generalization bounds for individual objectives extend to MOO with scalarization. Both studies rely on the Rademacher complexity of the hypothesis class to establish algorithm-independent generalization bounds, which are unaffected by the training process.
@@ 6.2. Generalization
(44) $\max_{i\in[m]}\ell_{{\mathcal{D}}_{i}}(h)\leq\min_{h^{*}\in{\mathcal{H}}}\max_{i\in[m]}\ell_{{\mathcal{D}}_{i}}(h^{*})+\epsilon,$
where $\ell_{{\mathcal{D}}_{i}}(h)$ is the loss of hypothesis $h$ on ${\mathcal{D}}_{i}$, $h^{*}$ is the optimal hypothesis and $\epsilon$ is the error tolerance. It is first formulated by Awasthi et al. (Awasthi et al., 2023) as an open problem in 2023 and has since been addressed by several works (Peng, 2024; Zhang et al., 2024c). Haghtalab et al. are the first to show the sample complexity lower bound of $\widetilde{\Omega}\left(\frac{v+m}{\epsilon^{2}}\right)$, where $v=\operatorname{VCdim}({\mathcal{H}})$ is the VC-dimension of the hypothesis class $\mathcal{H}$}. By using a boosting framework, Peng (Peng, 2024) gives an algorithm that achieves a sample complexity upper bound of $\widetilde{O}\left(\frac{v+m}{\epsilon^{2}}\cdot\left(\frac{m}{\epsilon}\right)^{o(1)}\right)$, which is nearly optimal. In a concurrent work, Zhang et al. (Zhang et al., 2024c) provide a variant of the hedge algorithm under an Empirical Risk Minimization (ERM) oracle access that achieves the optimal sample complexity bound of $\widetilde{O}\left(\frac{v+m}{\epsilon^{2}}\right)$.
In the online setting, Liu et al. (Liu et al., 2024c) provide an adaptive online mirror descent algorithm that achieves $\mathcal{O}\left(\frac{mv}{\sqrt{K}}\right)$ regret, where $K$ is the number of iterations. Using a plain online-to-batch conversion scheme, this algorithm leads to an $\mathcal{O}\left(\frac{m^{2}v^{2}}{\epsilon^{2}}\right)$ sample complexity, which still lags behind the optimal offline sample complexity of $\widetilde{O}\left(\frac{m+v}{\epsilon^{2}}\right)$ (Zhang et al., 2024c). It is interesting to see whether it is possible to use an online learner with proper online-to-batch conversion schemes that is able to match the optimal offline sample complexity.
MoDo (Chen et al., 2023a) introduces a different perspective by leveraging algorithm stability to derive algorithm-dependent generalization error bounds of gradient balancing algorithms.
7. Applications
7.1. Reinforcement Learning
Reinforcement Learning (RL) (Wang et al., 2024d) involves training agents to make sequential decisions by maximizing cumulative rewards. Multi-Objective Reinforcement Learning (MORL) (Felten et al., 2024) extends single-objective RL by utilizing a vector-valued reward function ${\bm{r}}(s,a):\mathcal{S}\times\mathcal{A}\mapsto\mathbb{R}^{m}$. MORL aims to learn a policy network $\pi_{\bm{\theta}}(s)$ with the following objective:
(45) $\min_{{\bm{\theta}}}{\bm{f}}({\bm{\theta}}):=\mathbb{E}_{a\sim\pi_{\bm{\theta}}(s)}\sum_{t=1}^{\infty}\gamma^{t}{\bm{r}}(s_{t},a_{t}),$